# Integration of TinyML and LargeML: A Survey of 6G and Beyond

### v1

@@ IV-B2 Advanced Federated Transfer Learning
Building upon the foundations of FTL, below are efficient methods for integrating TinyML with LargeML.
Federated Meta-Learning (FML)
@@
FML aims to achieve a higher level of generalization and adaptability by enabling models to learn new tasks rapidly through meta-knowledge acquired from diverse clients or domains [115]. Unlike standard FTL, which focuses on transferring knowledge across related tasks, FML trains models to adapt to previously unseen tasks. This “learning how to learn” capability improves robustness and broadens applicability, supporting both knowledge transfer and efficient task-specific adaptation.
@@
TinyReptile, an FML-based TinyML–LargeML system introduced in [116], draws on concepts from FL and online learning. The system enables distributed resource-constrained IoT devices to collaboratively learn a robust RNN initialization for rapid adaptation to new target devices. The study evaluates TinyReptile on classification tasks using sine-wave, Omniglot, and keyword spotting datasets and demonstrates a simple yet efficient implementation on Raspberry Pi 4 and Cortex-M4 MCU edge hardware, using parameter-based TL. TinyMetaFed, a comparable framework, is introduced in [117]. Evaluated on similar TinyML devices, TinyMetaFed incorporates a modified algorithm that reduces energy consumption and communication overhead, accelerates convergence, and stabilizes training. Building on the capabilities of TinyReptile and TinyMetaFed, the study in [118] proposes an advanced prototype, deployed on similar TinyML testbed devices. The framework employs a CNN backbone architecture for pre-trained models and applies the MLPerf Tiny Benchmark for on-device, personalized model adaptation. The prototype demonstrated its effectiveness in three real-world TinyML applications: handwritten character image classification, keyword spotting, and presence detection in smart buildings. PMFed, proposed in [119], is an FML framework that employs a pre-trained CNN and fine-tunes the model on a Raspberry Pi 3B for personalized adaptation on TinyML devices. The framework demonstrated efficient performance in three real-world, IoT-enabled health monitoring tasks: prompt defibrillation following detection of ventricular arrhythmia, atrial fibrillation detection, and human activity recognition.
Federated Foundation Models (FFMs) and Federated Fine-Tuning (FedFT) of Foundation Models (FMs)
@@
FFMs integrate the strengths of FMs and FL, enabling decentralized clients to collaboratively adapt pre-trained LargeML models without exchanging raw data. This approach combines the privacy advantages of FL with the TL benefits of TFMs [22]. FedFT of FMs allows each client to locally personalize a shared model while maintaining data confidentiality. To reduce computational overhead, PEFT techniques such as LoRA, adapter modules, and prompt tuning are widely applied [22].
@@
The study in [120] introduces FedPT, a federated proxy-tuning method for LLMs on resource-constrained edge devices. By relying solely on model predictions rather than parameter access, FedPT reconstructs full-sized LoRA modules for server-side aggregation, reducing computational and communication overhead while matching FedPT performance in instruction-following tasks. However, its reliance on homogeneous LoRA hampers performance under non-IID data distributions. To address this, HetLoRA [121] uses zero-padding and rank-constrained truncation to standardize LoRA modules, enhancing convergence and efficiency in heterogeneous systems for tasks like multi-session chat dialogue and text summarization. CAFF [122] further improves accuracy and fairness in language and vision classification tasks by combining tiny transformers with a layer-wise FedFT and NN selection strategy. In [123], the authors present FedD2P, an efficient FedFT framework for FMs targeting heterogeneous, resource-constrained IoT devices. FedD2P distills aggregated knowledge into a prompt generator, allowing frozen FMs to adapt efficiently to downstream tasks such as image classification, satellite imagery analysis, and texture recognition. The work in [124] proposes FwdLLM, an efficient FedFT method for LLMs on heterogeneous, resource-constrained edge devices. Instead of applying traditional BP to compute precise gradients, FwdLLM combines BP-free training with PEFT by injecting minor self-generated perturbations into model parameters. Compared to existing PEFT benchmarks, this approach reduces memory footprint and accelerates convergence.
@@
To further support model heterogeneity and computational efficiency, sparse architectures like MoE can be combined with PEFT. For example, FedFMSL [125] incorporates a two-stage MoE-based FedFT. The method first trains a global expert, then fine-tunes a local expert for improved personalization. To adapt to heterogeneous, resource-constrained edge environments, FedFMSL employs a sparsely-activated LoRA algorithm that progressively activates LoRA matrices based on historical performance tracked in a capability queue. This approach yields performance gains and resource savings in various image classification tasks, for example food recognition, land use, land cover, and human action detection. In [126], the authors introduce FedMoE, a sparsely-activated MoE operating with LoRA-based FedFT for heterogeneous, resource-constrained edge environments. FedMoE dynamically adjusts submodels using global expert recommendations and achieves fast convergence and improved performance in text classification, summarization, and reading comprehension tasks. Lastly, A3SMoE [48] introduces an efficient FedFT method for FMs tailored to heterogeneous, resource-restricted edge settings. A3SMoE extends the concept of data-conditional computation in sparsely-activated MoEs to resource-conditional computation, enabling clients to activate suitable experts based on resource availability. It outperforms prior LoRA-based techniques in instruction-tuning and complements existing heterogeneous LoRA methods.
@@ IV-C Split Learning
SL is an ML technique designed for collaborative model training under resource-constrained environments, enabling full model training on devices to improve learning efficiency [75]. SL is particularly suitable for integrating TinyML and LargeML within 6G and future network architectures. By partitioning the model across client devices and servers, SL allows TinyML devices to participate in training and inference tasks without becoming overwhelmed by the computational demands of LargeML while also preserving data privacy. Two main SL variants can be applied to TinyML–LargeML systems: vanilla SL (VSL) and parallel SL (PSL).
@@ IV-C1 Vanilla Split Learning
VSL is the foundational form of SL in which the training process proceeds sequentially fro client to client [127]. On the server side, either a standalone server architecture (e.g., two-tiered client-edge or two-tiered client-cloud) or a multi-level server collaboration model can be implemented.
[…]
@@
The integration of VSL-based TinyML with LargeML has enabled a diverse range of applications for tasks in resource-constrained environments. For example, a two-stage wireless channel adaptive split inference method [130] addresses memory and energy constraints on edge devices. The experimental results demonstrate enhanced privacy, improved inference accuracy, and dynamic adjustment of split points based on real-time variations in wireless channel gain. ARES [131] is an adaptive, resource-aware SL scheme designed for efficient model training in resource-constrained IoT systems. ARES mitigates the impact of stragglers by employing device-specific split layers, balances energy consumption with training time, and adapts to fluctuating network throughput and computing capacity. An adaptive SL algorithm based on Lyapunov theory is introduced in [132]. This approach dynamically selects split layers for resource-constrained devices and allocates server-side computing resources in wireless edge networks. Simulations demonstrate significant reductions in average training delay and energy consumption.
When multi-level server collaboration is employed, multiple cloud and edge servers jointly participate in VSL through a strategy known as multi-hop SL [75]. In this architecture, multiple hierarchical servers are organized in sequence, forming a serial multi-hop SL framework, as described in [133]. Here, the model is partitioned into multiple segments: (i) initial layers are deployed on TinyML devices; (ii) middle layers are distributed across multiple servers, each managing a distinct segment; (iii) final layers reside on a server that generates the output. This setup follows a similar structure to standalone server architectures but involves sequential inference and BP across multiple servers [134, 135, 133]. Specifically, the first client processes raw local data through its local model segment to generate smashed data before forwarding it to the first server. The server processes this data through its model segment and forwards the output to the next server, continuing through the chain until the final output is produced. During BP, the final server computes the loss, performs gradient updates, and propagates gradients backward through the server chain. Each intermediate server updates its own parameters and passes gradients to the preceding server or client. Clients then perform local BP and parameter updates in sequence, completing the training loop [134, 135]. In a system with $N$ TinyML devices and $K$ collaborating servers, each inner iteration requires $2NK$ inference and BP steps, highlighting the need for efficient algorithms to reduce latency.
@@ IV-C2 Parallel Split Learning
The sequential training process of VSL introduces substantial training latency and does not scale as the number of participating TinyML devices increases, a problem that will be exacerbated by the massive connectivity expected in 6G networks. The highly non-IID nature of client data can also degrade model performance in sequential training, since the model may become biased toward the data distribution of the previously trained client. To address these challenges, PSL enables multiple TinyML devices to train their local model segments concurrently [136, 137]. This parallelization reduces total training time, introduces “semi-scalability” relative to the number of participating devices, and partially mitigates the impact of non-IID data in 6G mediums.
@@
    Simultaneous client-side local FP: Each TinyML device ${\cal N}=\left\{{1,\ldots,N}\right\}$ trains on its local datasets ${{\cal D}_{n}}=\left\{{1,\ldots,{D_{n}}}\right\}$ using its model segment ${\bf{T}}_{n}$, generating smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the designated split layer.
@@
    FP of smashed data: Each device transmits its smashed data and corresponding labels $\left\{{{{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right),{\bf{y}}_{n}}\right\}$ to the server. The server receives these inputs concurrently from all devices.
@@
    BP of smashed data gradients: The computed gradients ¯$\nabla\bar{\cal L}({\bf{T}}_{n}({\bf{x}}_{n}))$ are returned to each respective TinyML device in parallel.
@@
$\mathop{\min}\limits_{{\bf{L}},{{\left\{{{{\bf{T}}_{n}}}\right\}}_{\forall n}}}\frac{{\sum\nolimits_{n\in{\cal N}}{\sum\nolimits_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{\bf{L}}\left({{{\bf{T}}_{n}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}\right)}}}}{{\sum\nolimits_{n\in{\cal N}}{{D_{n}}}}}.$ (4)
@@
Recent publications have applied PSL to TinyML–LargeML systems [140, 141, 142]. For example, the study in [140] proposes a privacy-sensitive PSL framework that supports SL training across multiple resource-constrained devices in parallel. The framework addresses overfitting caused by non-IID training orders and varying dataset sizes on resource-constrained devices while maintaining data privacy. Introduced in [141], AdaSplit represents an efficient scaling SL method for heterogeneous, resource-constrained systems. AdaSplit minimizes on-device computation and optimizes server-side processing and communication. It eliminates the need for server gradients at the client side, reduces payload size and interaction frequency to reduce communication overhead, and support updates to sparse model partitions on heterogeneous resource-constrained devices. The work in [142] formulates and solves a multiplayer bargaining problem to identify optimal split-layer strategies in PSL-based bidirectional integration systems. The approach balances computation and wireless transmission energy consumption, training time, and data privacy. The proposed framework also demonstrates robustness with non-IID datasets across resource-constrained devices.
In PSL-based TinyML–LargeML systems involving multi-level server collaboration, the architecture extends beyond simple client–edge or client–cloud configurations by incorporating multiple servers or hierarchical layers of servers that collaborate during the training process. Such setups require server orchestration and aggregation strategies, necessitating advanced PSL schemes to coordinate interactions between server layers and integrate TinyML with LargeML seamlessly. Detailed in Section IV-C3, this class of schemes is referred to as advanced PSL, where parallel-server PSL [143] and multi-hop PSL [144] serve as pioneering frameworks. These frameworks support the development of more sophisticated server orchestration and aggregation methods.
@@ IV-C3 Advanced Parallel Split Learning
EdgeSplit [138] is an advanced PSL variant designed to address the challenge of highly non-IID data distributed across TinyML devices. Non-IID data causes local models trained on individual TinyML devices to diverge, reducing generalization and degrading global model performance. EdgeSplit introduces a coordinated learning approach across diverse datasets while preserving the core benefits of PSL. Fig. 7 illustrates the inner-iteration training steps of EdgeSplit, which begins by a standard PSL procedure: 1 Simultaneous client-side local FP, 2 FP1 of smashed data, 3 server-side local FP and local BP, 4 BP1 of smashed data gradients, and 5 simultaneous client-side local BP. The subsequent steps introduce the key advancements of EdgeSplit over conventional PSL.
@@
    ${{\bf{T}}_{{\rm{FedAvg}}}}={{\sum\nolimits_{n\in{\cal N}}{{D_{n}}{{\bf{T}}_{n}}}}}\Big{/}{{\sum\nolimits_{n\in{\cal N}}{{D_{n}}}}}.$ (5)
@@
EdgeSplit is applied to a defined subset of TinyML devices. In the context of massive 6G connectivity, device heterogeneity is the expected norm, with differences in computing power, data distribution, and channel conditions being significant. In such scenarios, cluster-based PSL (CPSL) [145] offers a solution by building on the procedural structure of EdgeSplit. Instead of treating all devices as a uniform group, CPSL first clusters them according to shared characteristics (e.g., computing capabilities and channel conditions). These clusters contain devices with more manageable and stable data distributions, enabling more efficient training and improved model performance.
@@
Conventional PSL methods, such as EdgeSplit and CPSL, achieve scalability on the client side, whereas the server continues to carry a significant computational burden. EPSL [139] advances PSL by addressing this imbalance, targeting the issue of semi-scalability through a modified aggregation strategy. This strategy reduces server-side computational costs and training latency, enabling full scalability across both client and server tiers. The key steps of EPSL are illustrated in Fig. 8 and summarized below.
@@
    Simultaneous client-side local FP: Each TinyML device $n\in{\cal N}=\left\{{1,\ldots,N}\right\}$ independently trains on its local dataset ${{\cal D}_{n}}=\left\{{1,\ldots,{D_{n}}}\right\}$ using its local model segment, ${\bf{T}}_{n}$. The device generates smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the designated split layer. This step is executed in parallel across all devices.
@@
    FP of smashed data: Devices transmit their smashed data and corresponding labels, $\left\{{{{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right),{\bf{y}}_{n}}\right\}$, to the server. The server processes these multiple smashed data streams concurrently.
@@
    Client-side local BP: Each TinyML device uses the received gradients ${\bf{G}}_{a}$ and ${\bf{G}}_{u}$ to compute its local gradient ${\bf{G}}_{n}$ and update its model segment according to ${{\bf{T}}_{n}}\leftarrow{{\bf{T}}_{n}}-{\eta_{n}}{{\bf{G}}_{n}}$.
@@
Building on the PSL-based multi-level server collaboration framework introduced in Section IV-C2, we now examine parallel-server PSL and multi-hop PSL in detail. Parallel-server PSL [143] enables multiple parallel servers to process model segments concurrently for a set of resource-constrained clients. This setup introduces orchestration management and aggregation strategies that require the solution of joint optimization problems related to client-server assignments and scheduling decisions that minimize total training time. To address this NP-hard problem, we propose two scalable solutions: (i) a decomposition-based approach that exploits structural symmetry, and (ii) a lightweight alternative with low computational overhead. In contrast, multi-hop PSL [144] implements a pipelined, parallel multi-hop SL, extending the multi-hop SL approach discussed in Section IV-C1 with a parallel execution mechanism. Specifically, all participating TinyML devices simultaneously train on their local datasets using their respective model segments, generate smashed data at the split layer, and forward this data to the first server. The subsequent inference and backward paths between the servers proceed sequentially, following the same steps as in multi-hop SL. Until the BP process completes at the first server, all participating TinyML devices receive their corresponding smashed data gradients from this server, ensuring synchronization during the final aggregation phase. Unlike multi-hop SL, multi-hop PSL reduces the number of inference and backward path steps to $(2K+2)$, regardless of the value of $N$.
IV-D Federated Split Learning

### v2

@@ IV-B2 Advanced Federated Transfer Learning
Federated Meta-Learning (FML) aims to achieve a higher level of generalization and adaptability by enabling models to learn new tasks rapidly through meta-knowledge acquired from diverse clients or domains [106]. Unlike standard FTL, transferring knowledge across related tasks, FML trains models to adapt to previously unseen tasks. This “learning how to learn” capability improves robustness and broadens applicability, supporting both knowledge transfer and efficient task-specific adaptation.
@@
Several studies have explored FML for TinyML–LargeML integration. TinyReptile [141] combines FL and online learning to collaboratively train RNN initializations on resource-constrained IoT devices, enabling rapid adaptation to new targets. Evaluated on sine-wave, Omniglot, and keyword spotting datasets, it demonstrated efficient implementation on Raspberry Pi 4 and Cortex-M4 MCU using parameter-based TL. TinyMetaFed [140] extends this approach with an algorithm that reduces energy consumption and communication overhead, accelerates convergence, and stabilizes training on similar TinyML devices. Building on both TinyReptile and TinyMetaFed, the framework in [142] introduces an advanced prototype with a CNN backbone and MLPerf Tiny Benchmark, supporting personalized adaptation for handwritten character recognition, keyword spotting, and smart-building presence detection. Finally, PMFed [73] applies a pre-trained CNN fine-tuned on the Raspberry Pi 3B for personalized adaptation in IoT health monitoring, achieving efficient performance in defibrillation support, atrial fibrillation detection, and human activity recognition.
@@
On the other hand, Federated Foundation Models (FFMs) integrate the strengths of Foundation Models (FMs) and FL, enabling decentralized clients to collaboratively adapt pre-trained LargeML models without exchanging raw data. This approach combines the privacy advantages of FL with the TL benefits of TFMs [139]. Federated Fine-Tuning (FedFT) of FMs allows each client to locally personalize a shared model while maintaining data confidentiality. To reduce computational overhead, PEFT techniques such as LoRA, adapter modules, and prompt tuning are widely applied [139].
@@
Recent work has advanced FFMs and FedFT for FMs on resource-constrained edge devices. FedPT [50] introduces a federated proxy-tuning method, reconstructing full-sized LoRA modules from predictions for server-side aggregation, reducing computational and communication overhead but struggling under non-IID data. To address heterogeneity, HetLoRA [29] standardizes LoRA modules via zero-padding and rank truncation, improving convergence in tasks such as multi-session chat dialogue and text summarization. CAFF [130] further improves accuracy and fairness in language and vision classification tasks by combining tiny transformers with a layer-wise FedFT and NN selection strategy. FedD2P [11] distills aggregated knowledge into a prompt generator, enabling frozen FMs to adapt efficiently to downstream vision tasks (e.g., image classification or satellite imagery analysis). Moreover, FwdLLM [192] reduces footprint and accelerates convergence by using BP-free training with PEFT and injecting minor self-generated perturbations into model parameters.
@@
To further improve efficiency in heterogeneous environments, several studies integrate sparse architectures like MoE with PEFT. For instance, FedFMSL [190] incorporates a two-stage MoE-based FedFT, training global experts before fine-tuning local experts for improved personalization, with sparsely-activated LoRA matrices guided by historical performance tracked in a capability queue. This approach yields performance gains and resource savings in various image classification tasks, such as food recognition, land use, land cover, and human action detection. FedMoE [114] extends FedFMSL by dynamically activating submodels via global expert recommendations, achieving faster convergence in NLP tasks such as text classification, summarization, and reading comprehension. Finally, A3SMoE [166] introduces resource-conditional computation, activating sparsely-activated MoEs based on client resource availability, surpassing prior LoRA-based techniques in instruction-tuning and complements existing heterogeneous LoRA methods.
@@ IV-C Split Learning
SL is an ML paradigm for collaborative model training in resource-constrained environments, allowing devices to jointly train a complete model while enhancing learning efficiency (see [102] for more details). SL is particularly suitable for integrating TinyML and LargeML within 6G and future network architectures. By partitioning the model across client devices and servers, SL allows TinyML devices to participate in training and inference tasks without becoming overwhelmed by the computational demands of LargeML while also preserving data privacy. Two main SL variants can be applied to TinyML–LargeML systems: vanilla SL (VSL) and parallel SL (PSL).
@@ IV-C1 Vanilla Split Learning
VSL is the foundational form of SL in which the training process proceeds sequentially between clients and a coordinating server [175]. The server can follow a standalone architecture (e.g., client-edge or client-cloud) or a multi-level collaborative model.
Fig. 6(a) illustrates a VSL-based integration of TinyML–LargeML, in which a standalone cloud or edge server collaborates with multiple TinyML devices during training. Typically, a smaller sub-model ${\bf{T}}_{n}$ resides on each TinyML device, and a larger sub-model $\bf{L}$ is hosted on the server [175]. Each client $n\in{\cal N}$ trains on its local dataset ${\cal D}_{n}=\left\{{{\bf{x}}_{j}^{n},{\bf{y}}_{j}^{n}}\right\}_{j=1}^{{|{\cal D}_{n}|}}$. Together, they form a temporary global model ${{\bf{G}}_{n}}=\left\{{{{\bf{T}}_{n}};{\bf{L}}}\right\}$. Training proceeds over $R$ outer rounds, where clients interact with the server in a round-robin fashion. For each client, $T$ inner iterations are executed before moving to the next. During each inner iteration, it particularly proceeds:
  • Step $(n.1)$: Client $n$ performs forward propagation (FP) through its local model ${\bf{T}}_{n}$, up to the designated split (or cut) layer. The objective is to minimize its local loss function $\bar{\cal L}\left({{{\bf{G}}_{n}}|{{\cal D}_{n}}}\right)\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\frac{1}{{{|{\cal D}_{n}|}}}\sum\nolimits_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{{\bf{G}}_{n}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}$. This produces intermediate feature representations, called smashed data, ${{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right)$}.
  • Step $(n.2)$: The smashed data and labels, $\left\{{{{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right),{\bf{y}}_{n}}\right\}$, are transmitted to the server. Data may be sent in full batches or mini-batches depending on optimal hyperparameters or task-specific requirements [91].
  • Step $(n.3)$: The server continues FP using $\bf{L}$, computes the gradients of the loss function $\bar{\cal L}\left({{{\bf{G}}_{n}}|{{\cal D}_{n}}}\right)$ to generate $\nabla{\bar{\cal L}}({\bf{L}})$ and $\nabla{\bar{\cal L}}({\bf{T}}_{n}({\bf{x}}_{n}))$. Then, it performs server-side local backward propagation (BP) by updating ${\bf{L}}\leftarrow{\bf{L}}-\eta_{S}\nabla{\bar{\cal L}}({\bf{L}})$, where $\eta_{S}$ is the server learning rate.
  • Step $(n.4)$: The server sends the smashed data gradients $\nabla{\bar{\cal L}}({\bf{T}}_{n}({\bf{x}}_{n}))$ back to client $n$.
  • Step $(n.5)$: Client $n$ updates its local sub-model as ${\bf{T}}_{n}\leftarrow{\bf{T}}_{n}-\eta_{n}\nabla{\bar{\cal L}}({\bf{T}}_{n}({\bf{x}}_{n}))$, where $\eta_{n}$ is the client learning rate. For consistency, the terms $\eta_{S}$ and $\eta_{n}$ are reused with the same definitions in subsequent sections.
This exchange completes one inner iteration. During the outer iterations, the aim is to minimize the global loss function $\mathop{\min}\limits_{{{\{\bf{G}}_{n}}\}_{\forall n}}{{\sum\nolimits_{n\in{\cal N}}{{|{\cal D}_{n}|}\bar{\cal L}\left({{{\bf{G}}_{n}}|{{\cal D}_{n}}}\right)}}}\big/{{\sum\nolimits_{n\in{\cal N}}{{|{\cal D}_{n}|}}}}$. After convergence, TinyML devices can employ the global model $\bf{G}$} for inference on a designated target task.
@@
The integration of VSL-based TinyML with LargeML has enabled a diverse range of applications for tasks in resource-constrained environments. Specifically, a two-stage wireless channel adaptive split inference method [88] addresses edge device memory and energy limits by dynamically adjusting split points based on real-time channel gain, improving privacy, inference accuracy, and adaptability. Similarly, ARES [148], an adaptive resource-aware SL scheme, mitigates stragglers through device-specific split layers, balancing energy use, training time, and throughput variability. An adaptive SL algorithm based on Lyapunov theory [95] further reduces training delay and energy consumption by dynamically selecting split layers and allocating server-side resources in edge networks.
When extended to multi-level server collaboration, VSL can be deployed via multi-hop SL [102], where cloud and edge servers form a hierarchical chain. In this setup, initial layers run on TinyML devices, intermediate layers on distributed servers, and final layers on a central server [181]. During training, inference proceeds forward through the network while BP sends gradients backward to update each node’s parameters [160, 22]. With $N$ TinyML devices and $K$ servers, each inner iteration requires $2NK$ inference and BP steps, highlighting the need for latency‑efficient algorithms.
@@ IV-C2 Parallel Split Learning
The sequential training process of VSL leads to significant training latency and struggles to scale with more TinyML devices, particularly with the anticipated connectivity in 6G networks. Additionally, the non-IID nature of client data can bias model performance. PSL addresses these issues by allowing multiple TinyML devices to train their models concurrently, which reduces training time [111], offers “semi-scalability” with more devices, and helps mitigate the effects of non-IID data in 6G environments.
@@
    Simultaneous client-side local FP: Each TinyML device $n\in{\cal N}$ trains on its local dataset ${\cal D}_{n}$} using its model segment ${\bf{T}}_{n}$, generating smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the designated split layer.
@@
    FP of smashed data: Each device transmits its smashed data and labels, $\left\{{{{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right),{\bf{y}}_{n}}\right\}$, to the server. The server receives these inputs concurrently from all devices.
@@
    BP of smashed data gradients: The computed gradients $\nabla\bar{\cal L}({\bf{T}}_{n}({\bf{x}}_{n}))$ are returned to each respective TinyML device in parallel.
@@
$\mathop{\min}\limits_{{\bf{L}},{{\left\{{{{\bf{T}}_{n}}}\right\}}_{\forall n}}}{\sum_{n\in{\cal N}}{\sum_{j\in{{\cal D}_{n}}}{{\cal L}\left({{\bf{y}}_{j}^{n},{\bf{L}}\left({{{\bf{T}}_{n}}\left({{\bf{x}}_{j}^{n}}\right)}\right)}\right)}}}\frac{1}{{\sum\nolimits_{n\in{\cal N}}{{|{\cal D}_{n}|}}}}.$ (3)
@@
Recent studies have applied PSL to TinyML–LargeML systems [72, 30, 79]. A privacy-sensitive PSL framework in [72] enables parallel SL training across multiple resource-constrained devices, mitigating overfitting from non-IID training orders and varying dataset sizes on resource-constrained devices while preserving privacy. AdaSplit [30] enhances scalability in heterogeneous, resource-constrained systems by minimizing on-device computation, optimizing server-side processing, and reducing via sparse model updates without requiring server gradients on clients. In [79], PSL-based bidirectional integration is modeled as a multiplayer bargaining problem to identify optimal split-layer strategies, balancing computation, transmission energy, training time, and data privacy, while showing robustness on non-IID datasets.
In PSL-based TinyML–LargeML systems involving multi-level server collaboration, the architecture extends beyond simple client-edge or client-cloud configurations by incorporating multiple servers or hierarchical layers of servers that collaborate during the training process. Such setups require server orchestration and aggregation strategies, necessitating advanced PSL schemes to coordinate interactions between server layers and integrate TinyML with LargeML seamlessly. Detailed in Section IV-C3, this class of schemes is referred to as advanced PSL, where parallel-server PSL [162] and multi-hop PSL [161] serve as pioneering frameworks. These frameworks support the development of more sophisticated server orchestration and aggregation methods.
@@ IV-C3 Advanced Parallel Split Learning
EdgeSplit [204] is an advanced PSL variant designed to address the challenge of highly non-IID data distributed across TinyML devices. Non-IID data causes local models trained on individual TinyML devices to diverge, reducing generalization and degrading global model performance. EdgeSplit introduces a coordinated learning approach across diverse datasets while preserving the core benefits of PSL. Fig. 7 illustrates the inner-iteration training steps of EdgeSplit, which begins by a standard PSL procedure: 1 Simultaneous client-side local FP, 2 FP1 of smashed data, 3 server-side local FP and local BP, 4 BP1 of smashed data gradients, and 5 simultaneous client-side local BP. The subsequent steps introduce the key advances of EdgeSplit over conventional PSL.
@@
    ${{\bf{T}}_{{\rm{FedAvg}}}}={{\sum\nolimits_{n\in{\cal N}}{{|{\cal D}_{n}|}{{\bf{T}}_{n}}}}}\Big/{{\sum\nolimits_{n\in{\cal N}}{{|{\cal D}_{n}|}}}}.$ (4)
@@
EdgeSplit is used for selected TinyML devices. In the context of widespread 6G connectivity, device heterogeneity is common, leading to variations in computing power, data distribution, and channel conditions. Cluster-based PSL (CPSL) [191] addresses this by clustering devices based on shared characteristics like computing capabilities. This results in clusters with stable data distributions, enabling more efficient training and enhanced model performance.
@@
Conventional PSL methods like EdgeSplit and CPSL scale well on the client side but still leave heavy computation on the server. EPSL [101] addresses this imbalance with a revised aggregation strategy that lowers server‑side cost and training latency, achieving full end‑to‑end scalability. Its main steps are shown in Fig. 8 and summarized below.
@@
    Simultaneous client-side local FP: Each TinyML device $n\in{\cal N}$ independently trains on its local dataset ${\cal D}_{n}$} using its local model segment, ${\bf{T}}_{n}$. The device generates smashed data ${\bf{T}}_{n}({\bf{x}}_{n})$ at the designated split layer. This step is executed in parallel across all devices.
@@
    FP of smashed data: Devices transmit their smashed data and labels, $\left\{{{{\bf{T}}_{n}}\left({{\bf{x}}_{n}}\right),{\bf{y}}_{n}}\right\}$, to the server. The server processes these multiple smashed data streams concurrently.
@@
    Client-side local BP: Each TinyML device uses the received gradients ${\bf{G}}_{a}$ and ${\bf{G}}_{u}$ to compute its local gradient ${\bf{G}}_{n}$ and update its model segment ${{\bf{T}}_{n}}\leftarrow{{\bf{T}}_{n}}-{\eta_{n}}{{\bf{G}}_{n}}$.
@@
From the PSL-based multi-level server collaboration framework introduced in Section IV-C2, two extensions have been proposed: parallel-server PSL and multi-hop PSL. In parallel-server PSL [162], multiple servers concurrently process model segments for resource-constrained clients, requiring orchestration management and joint optimization of client-server assignments and scheduling to minimize training time. Due to its NP-hard nature, two scalable solutions are suggested: (i) a decomposition-based method utilizing structural symmetry, and (ii) a lightweight alternative with low computational overhead. Conversely, multi-hop PSL [161] enhances traditional multi-hop SL with pipelined, parallel execution. Each locally trained TinyML device generates smashed data at a split layer and sends it to the first server, where inference and backpropagation occur sequentially, akin to multi-hop SL in Section IV-C1. After backpropagation, devices receive synchronized gradient updates for aggregation. Unlike multi-hop SL, multi-hop PSL reduces inference and backward steps to $(2K+2)$, independent of the number of devices $N$.
@@
IV-D Federated Split Learning (FSL)
PSL parallelizes client‑side model training (see Section IV-C2) while handling highly non‑IID client data, effectively blending the advantages of split learning and federated learning [159]. FSL differs from PSL, EdgeSplit, and EPSL by distributing responsibilities across two cooperating servers: a main cloud server and a Fed server that runs FedAvg. In FSL, the Fed server aggregates updated local models from clients using FL-based techniques (see [170] for more details).